OpenStudy (caitlinnr14):

PLEASE HELPPP WILL GIVE MEDAL! Find the product of (3x − 4)(2x^2 + 2x − 1).

OpenStudy (***[isuru]***):

r u asking for the expanded form ?

OpenStudy (caitlinnr14):

no like it ..its hard to explain ...hold on one sec.

OpenStudy (***[isuru]***):

for example (x-1)(x-2) = x^2 -3x +3 something like that ?

OpenStudy (caitlinnr14):

Here are the answers anything after the x is an exponent 6x3 + 2x2 − 5x + 4 6x3 + 14x2 − 11x + 4 6x3 − 14x2 − 5x + 4 6x3 − 2x2 − 11x + 4

OpenStudy (caitlinnr14):

multiplication of polynomials

OpenStudy (***[isuru]***):

(3x − 4)(2x^2 + 2x − 1) =3x(2x^2 + 2x − 1) - 4(2x^2 + 2x − 1) = 6x^3 + 6x^2 -3x -8x^2 - -8x +4 = 6x^3 -2x^2 -11x +4 I think this is wt u were looking for and hope this will help u out

OpenStudy (caitlinnr14):

ok thankyou btw thx for fanning me. did i fan u yet?

OpenStudy (***[isuru]***):

u r welcome and u haven't fanned me yet ;-(

OpenStudy (caitlinnr14):

oh sry i will lol

OpenStudy (***[isuru]***):

thanks ^^